OpenStudy (anonymous):

give eg.. 1. An unbounded sequence that has a convergent subsequence 2. An unbounded sequence that has no convergent subsequence 3. A null sequence (an) such that the series ¡Æ an does not converge 4. A sequence which is not a Cauchy sequence but has the property that for every ¦Å > 0 and every N > 0 there exists n > N and m > 2n such that mod(an ¨C am) ¡Ü ¦Å 5. Two sequences (an) and (bn) such that the sequence (cn) defined by cn = an + bn converges to 1 but neither (an) nor (bn) converge. 6. A sequence (an) that tends to +ve infinity but is neither increasing nor eventually inc

OpenStudy (anonymous):

here the whole question comes Give examples of: 1. An unbounded sequence that has a convergent subsequence 2. An unbounded sequence that has no convergent subsequence 3. A null sequence (an) such that the series ¡Æ an does not converge 4. A sequence which is not a Cauchy sequence but has the property that for every ¦Å > 0 and every N > 0 there exists n > N and m > 2n such that mod(an ¨C am) ¡Ü ¦Å 5. Two sequences (an) and (bn) such that the sequence (cn) defined by cn = an + bn converges to 1 but neither (an) nor (bn) converge. 6. A sequence (an) that tends to +ve infinity but is neither increasing nor eventually increasing. 7. A bounded sequence (an) such that (an+1/an) tends to 1 but (an) doesn¡¯t converge. 8. A bounded sequence which doesn¡¯t converge. 9. A series ¡Æ an which converges but such that ¡Æ 3¡Ìan diverges. 10. A series ¡Æ an with sequence of partial sums (sn) such that sn = an for all natural numbers n. 11. A sequence (an) of non-zero terms such that (an+2/an) tends to 1 but (an) does not converge. 12. A sequence (an) of non-zero terms which is bounded above but such that the sequence (1/an) is not bounded below. 13. A convergent series ¡Æ an with all terms non-zero and ¡Æ an ¡Ù 7 which can be rearranged so that its sum is 7. 14. A series ¡Æ an which converges but such that ¡Æ (-1)^n+1.an does not converge. 15. A sequence (an) such that, for all natural numbers m, an = m for infinitely many natural numbers n. 16. A Cauchy sequence 17. A sequence of the form (x^n + y^n)^(1/n) which converges to 6, as n tends to infinity. 18. Two series ¡Æ an and ¡Æ bn which diverge to infinity but where ¡Æ an.bn converges. 19. A sequence (an) which satisfies the following but does not diverge to infinity: for all C > 0, there exists N so that there exists n > N with an > C 20. A sequence (an) with (1/an) tending to 0 but (an) not tending to infinity. 21. A non-monotonic, unbounded sequence. 22. A sequence (an) which satisfies the following but does not converge to 1: for all ¦Å > 0, for all N there exists n > N with mod(an ¨C 1) < ¦Å 23. A number which has only one decimal expansion 24. A sequence of irrationals which has a subsequence converging to 2 and a subsequence converging to 3. 25. A sequence (an) with (an+1/an) tending to 1 but which does not converge. 26. A series ¡Æ an, with non-zero terms, which can be rearranged so that its sum is 3.5. 27. A series ¡Æ an, with (an) tending to 0 and with (sn) tending to infinity , where (sn) is the sequence of partial sums. 28. A sequence (an) which tends to infinity but has an+1 < an for infinitely many n.

OpenStudy (anonymous):

1. An unbounded sequence that has a convergent subsequence \[ a_{2n}=2n\quad a_{2n+1}=\frac 1{n+1}\] 2. An unbounded sequence that has no convergent subsequence \[ a_{n}=n\] 3. A null sequence (an) such that the series ¡Æ an does not converge It is not clear what you want(unreadable characters) 4.A sequence which is not a Cauchy sequence but has the property that for every ¦Å > 0 and every N > 0 there exists n > N and m > 2n such that mod(an ¨C am) ¡Ü ¦Å It is not clear what you want (unreadable characters) 5. Two sequences (an) and (bn) such that the sequence (cn) defined by cn = an + bn converges to 1 but neither (an) nor (bn) converge. \[ a_{n}=n\quad b_n=-n+1 \] 6. A sequence (an) that tends to +ve infinity but is neither increasing nor eventually increasing. \[ a_{n}=\frac n 2 +\sin(n) \] 7. A bounded sequence (an) such that (an+1/an) tends to 1 but (an) doesn¡¯t converge. \[ a_{n}=n \] 8. A bounded sequence which doesn¡¯t converge. \[ a_{n}=sin(n) \] 9. A series ¡Æ an which converges but such that ¡Æ 3¡Ìan diverges. It is not clear what you want(unreadable characters) 10. A series ¡Æ an with sequence of partial sums (sn) such that sn = an for all natural numbers n. \[ a_{n}=0 \] 11. A sequence (an) of non-zero terms such that (an+2/an) tends to 1 but (an) does not converge. \[ a_{n}=n \] 12. A sequence (an) of non-zero terms which is bounded above but such that the sequence (1/an) is not bounded below. \[ a_{n}=\sin(n) \] 13. A convergent series ¡Æ an with all terms non-zero and ¡Æ an ¡Ù 7 which can be rearranged so that its sum is 7. It is not clear what you want(unreadable characters) 14. A series ¡Æ an which converges but such that ¡Æ (-1)^n+1.an does not converge. \[ a_n =(-1)^{n+1} \frac 1 n\] 15. A sequence (an) such that, for all natural numbers m, an = m for infinitely many natural numbers n. \[ a_{n,m}= n \text { for all n and m} \] It is an infinite matrix whose nth row is eqaul to n 16. A Cauchy sequence \[ a_n =\frac 1 n\] 17. A sequence of the form (x^n + y^n)^(1/n) which converges to 6, as n tends to infinity. \[ x=0 \quad y= 6 \quad (x^n + y^n)^{1/n} = 6 \] for all n. 18. Two series ¡Æ an and ¡Æ bn which diverge to infinity but where ¡Æ an.bn converges. \[ a_n = b_n = \frac 1 n \] 19. A sequence (an) which satisfies the following but does not diverge to infinity: for all C > 0, there exists N so that there exists n > N with an > C \[ a_{2n}=0\quad a_{2n+1}=2n+1\] 20. A sequence (an) with (1/an) tending to 0 but (an) not tending to infinity. \[ a_n = (-1)^n n \] This sequence does not tend to [\ +\infty\] You can do the rest.